# Critical Nodes Identification in Complex Networks: A Survey

### v1

@@ 8.2 Dynamical Sensitivity
Liu et al. liu2016locating has proposed the dynamic sensitive centrality (DSC) by integrating topological features and dynamic characteristics, considering a discrete time spreading model in which an infected node will infect its neighbors with a spreading rate $\beta$ and recover with a recovery rate $\mu$. Let $x(t)(t>0)$ denote the approximate value of the cumulative probability of nodes being activated between time step 1 and $t$, then $x(t)-x(t-1)(t>1)$ is approximated as the probability of a node being infected at time step $t$. If $v_{i}$ is the unique initial infected node, then $x_{i}(0)=1$ and $x_{j\neq i}(0)=0$. In the first time step, $x(1)=\beta Ax(0)$, and for $t>1$: $x(t)-x(t-1)=\beta A[\beta A+(1-\mu)I]^{t-1}x(0)$, where $v_{i}$ is the identity matrix. Let $H=\beta A+(1-\mu)I$, then the cumulative probability of nodes being infected between time step 1 and $t$ can be approximated as
@@
Gravity Formula RN14 ,  maji2020influential ,  RN16 ,  RN24 ,  liu2020gmm ,  RN13 ,  RN20 ,  fei2018identifying ,  RN28 ,  RN26 ,  yan2020identifying ,  wang2022identifying20 ,  yang2023aigcrank17 ,  zhu2023identifying12 ,  liu2023entropy12 ,  zhao2022random21 ,  zhao2023estimating3 ,  RN25 ,  curado2023novel ,  yang2023aogc ,  chen2023identification ,  meng2025improved ,  xu2024cagm ,  li2023improved123 ,  lv2024improved ,  chi2024measuring $+$ Comprehensive consideration of multiple factors, flexibility in handling various network structures. -- Potential complexity in parameter tuning, reliance on accurate index weightings.
Topsis and Entropy-Weight Multiple Metrics comin2011identifying ,  de2014role ,  hu2015multi ,  hu2015identifying ,  bucur2020top ,  wei2022identifying24 ,  an2024novel ,  wu2024hunting ,  cao2024dynamic ,  kopsidas2022identification ,  wang2024multi ,  lei2024weighted ,  ullah2024leveraging ,  zhang2024critical ,  lee2024identifying ,  esfandiari2025collaborative ,  chen2025sfimco ,  zhang2025novel ,  mo2015evidential ,  xu2025lgp ,  sheikhahmadi2017identification12 ,  wang2017ranking ,  sheikhahmadi2017identification ,  yang2018ranking ,  ullah2021identification ,  ullah2021identifying12 ,  hu2022exploring209 ,  wang2022influential ,  mukhtar2023integrating ,  yang2023identifying23 ,  qiu2024key ,  du2014new ,  liu2015node ,  hu2016modified ,  li2023identifying23 ,  dong2022cpr34 ,  chen2021effects ,  ishfaq2022identifying45 ,  yang2018multi ,  vega2019multi ,  lu2020node ,  zhang2022multi23 ,  ju2021novel23 $+$ Holistic analysis by integrating multiple evaluation criteria, increases accuracy in identifying critical nodes, adaptability to different network types. -- Difficulties in balancing and weighting diverse criteria effectively.
Information propagation method macdonald2012spreaders ,  borgatti2006identifying ,  zhuge2010topological ,  aral2012identifying ,  li2014efficient ,  lawyer2015understanding ,  chen2015critical ,  robinaugh2016identifying ,  bozorgi2016incim ,  holme2017three ,  yin2023identifying47 ,  he2019tifim ,  tulu2019vital ,  zhong2021information ,  li2021estimating ,  gong2021probability ,  wang2023hgim ,  mohammadi2024improved ,  wang2024identifying ,  xu2023novel ,  ai2022identifying10 ,  zareie2020identification ,  chen2020influential ,  fink2023centrality ,  sun2023finding ,  ullah2023lss ,  corsin2025evidence ,  liu2016locating ,  mo2019identifying ,  du2020identifying $+$ Effective identification of nodes crucial for information flow, applicability to diverse network types. -- Potential computational intensity, reliance on accurate modeling of propagation processes.
@@ 11 Summary and Discussion
Closeness Centrality CC freeman2002centrality ,EFF latora2001efficient ,GLR salavati2019ranking ,GLS sheng2020identifying
@@
Maximization of Connected Component CC-CNDP Boginski0007 , CC-CNP arulselvan2011cardinality ,BPT li2011finding ,MOCNDP ventresca2015experimental ,DSFLA tang2020discrete

### v2

@@ 8.2 Dynamical Sensitivity
Liu et al. liu2016locating has proposed the dynamic sensitive centrality (DSC) by integrating topological features and dynamic characteristics, considering a discrete time spreading model in which an infected node will infect its neighbors with a spreading rate $\beta$ and recover with a recovery rate $\mu$. Let $x(t)(t>0)$ denote the approximate value of the cumulative probability of nodes being activated between time step 1 and $t$, then x(t)x(t1)(t>1)x(t)-x(t-1)(t>1) is approximated as the probability of a node being infected at time step $t$. If $v_{i}$ is the unique initial infected node, then $x_{i}(0)=1$ and $x_{j\neq i}(0)=0$. In the first time step, $x(1)=\beta Ax(0)$, and for $t>1$: $x(t)-x(t-1)=\beta A[\beta A+(1-\mu)I]^{t-1}x(0)$, where $v_{i}$ is the identity matrix. Let $H=\beta A+(1-\mu)I$, then the cumulative probability of nodes being infected between time step 1 and $t$ can be approximated as
@@
Gravity Formula RN14 ,  maji2020influential ,  RN16 ,  RN24 ,  liu2020gmm ,  RN13 ,  RN20 ,  fei2018identifying ,  RN28 ,  RN26 ,  yan2020identifying ,  wang2022identifying20 ,  yang2023aigcrank17 ,  zhu2023identifying12 ,  liu2023entropy12 ,  zhao2022random21 ,  zhao2023estimating3 ,  RN25 ,  curado2023novel ,  yang2023aogc ,  chen2023identification ,  meng2025improved ,  xu2024cagm ,  li2023improved123 ,  lv2024improved ,  chi2024measuring ++ Comprehensive consideration of multiple factors, flexibility in handling various network structures. - Potential complexity in parameter tuning, reliance on accurate index weightings.
Topsis and Entropy-Weight Multiple Metrics comin2011identifying ,  de2014role ,  hu2015multi ,  hu2015identifying ,  bucur2020top ,  wei2022identifying24 ,  an2024novel ,  wu2024hunting ,  cao2024dynamic ,  kopsidas2022identification ,  wang2024multi ,  lei2024weighted ,  ullah2024leveraging ,  zhang2024critical ,  lee2024identifying ,  esfandiari2025collaborative ,  chen2025sfimco ,  zhang2025novel ,  mo2015evidential ,  xu2025lgp ,  sheikhahmadi2017identification12 ,  wang2017ranking ,  sheikhahmadi2017identification ,  yang2018ranking ,  ullah2021identification ,  ullah2021identifying12 ,  hu2022exploring209 ,  wang2022influential ,  mukhtar2023integrating ,  yang2023identifying23 ,  qiu2024key ,  du2014new ,  liu2015node ,  hu2016modified ,  li2023identifying23 ,  dong2022cpr34 ,  chen2021effects ,  ishfaq2022identifying45 ,  yang2018multi ,  vega2019multi ,  lu2020node ,  zhang2022multi23 ,  ju2021novel23 ++ Holistic analysis by integrating multiple evaluation criteria, increases accuracy in identifying critical nodes, adaptability to different network types. - Difficulties in balancing and weighting diverse criteria effectively.
Information propagation method macdonald2012spreaders ,  borgatti2006identifying ,  zhuge2010topological ,  aral2012identifying ,  li2014efficient ,  lawyer2015understanding ,  chen2015critical ,  robinaugh2016identifying ,  bozorgi2016incim ,  holme2017three ,  yin2023identifying47 ,  he2019tifim ,  tulu2019vital ,  zhong2021information ,  li2021estimating ,  gong2021probability ,  wang2023hgim ,  mohammadi2024improved ,  wang2024identifying ,  xu2023novel ,  ai2022identifying10 ,  zareie2020identification ,  chen2020influential ,  fink2023centrality ,  sun2023finding ,  ullah2023lss ,  corsin2025evidence ,  liu2016locating ,  mo2019identifying ,  du2020identifying ++ Effective identification of nodes crucial for information flow, applicability to diverse network types. - Potential computational intensity, reliance on accurate modeling of propagation processes.
@@ 11 Summary and Discussion
Closeness Centrality $C$ freeman2002centrality ,EFF latora2001efficient ,GLR salavati2019ranking ,GLS sheng2020identifying
@@
Maximization of Connected Component $C$-CNDP Boginski0007 , $C$-CNP arulselvan2011cardinality ,BPT li2011finding ,MOCNDP ventresca2015experimental ,DSFLA tang2020discrete